Please write clearly
balanced equation and find the molar ratio
1.O2 + N2 =N2O
2.P+O2=P2O5
3.Hz+Al=AlH3
4.C+H2=C2H6
5.P+H2=PH3
write the balanced equation and Molar ratio
1. reaction between fluorine and zinc to form zinc fluoride
2. nitrogen and oxygen to form nitrogen dioxide
3. oxygen and sodium to form sodium oxide
4.boron and nitrogen form boron nitride
5.lithium and bromine to form lithium bromide
Answer #1

1.02 + N2 = N20 Balanced Equation: 02 + 2N2 = 2N20 Molar ratio: 02: N2 = 1:2, 02: N2O=1:2, N2 : N20 = 1:1 . 2.P+02=P205 Balan4.C+H2=C2H6 Balanced Equation: 2C + 3H2 = C2H6 Molar ratio: C:H2 = 2:3, C : C2H6 = 2 : 1, H2 : C2H6 = 3 : 1 5. P+H2= PHz Bala1. Reaction between fluorine and zinc to form zinc fluoride Balanced Equation: F2 + Zn= ZnF2 Molar ratio: F2: Zn = 1:1, F2: Z4. Boron and nitrogen form boron nitride Balanced Equation: 2B + N2 = 2BN Molar ratio: B: N2 = 2 : 1, B: BN = 1:1, N2 : BN =